# Four-Plane Factorized Video Autoencoders

### v1

@@ 4 Experiments
4.1 Class-conditional generation and future frames prediction
We introduce the details of the class-conditional generation and frame prediction tasks together since their experimental settings share some similarities.
Datasets.
The class-conditional generation experiment is performed on the UCF-101 dataset [44] which contains 9,537 videos across 101 categories, covering a wide range of actions in total. The future frames prediction experiment is performed on the Kinetics-600 [6] dataset, consisting of a nearly $400,000$ video clips sourced from YouTube, containing around 600 action classes that cover a wide range of human activities.
Autoencoder pre-training.
Following W.A.L.T. [20], we pre-train the factorized autoencoder on the Kinetics-600 dataset for both the class-conditional generation and frame extrapolation tasks.
Diffusion model training.
For the class-conditional generation task, we train the diffusion model to generate four latent feature planes that decode to $128\times 128$ resolution videos with 17 frames, conditioned on the class label. This setup aligns with the approach used in W.A.L.T. [20] to ensure a fair comparison.
For the frame prediction task, we condition on the first spatial plane, $P_{xy}^{1}$, and generate the remaining three planes. Leveraging a causal encoder, as in MAGVIT-v2, enables this conditional generation to parallel the setup in W.A.L.T.’s frame prediction experiments, where the model is conditioned on two latent frames.
4.1.1 Metric
To evaluate the quality of generated videos, we use the Fréchet Video Distance (FVD) [45] as our primary metric. FVD measures the similarity between the distributions of generated and real videos, assessing both spatial realism and temporal coherence. Formally, FVD is defined as:
$\text{FVD}=\|\mu_{\text{real}}-\mu_{\text{gen}}\|^{2}+\operatorname{Tr}(\Sigma_{\text{real}}+\Sigma_{\text{gen}}-2(\Sigma_{\text{real}}\Sigma_{\text{gen}})^{1/2}),$
where $\mu_{\text{real}},\Sigma_{\text{real}}$ and $\mu_{\text{gen}},\Sigma_{\text{gen}}$ are the mean and covariance matrices of feature embeddings for real and generated
@@ 4.1.2 Analysis
As shown in  Table 1, our diffusion model, trained on the proposed factorized latent representation, outperforms most prior works and performs comparably to W.A.L.T. [20]. Additionally, our method is nearly 2×2$\times$2 × faster speeds for both training and inference compared to W.A.L.T. Under identical training architecture and computational resources, our model processes each training iteration in just $380$ milliseconds, compared to $750$ milliseconds for W.A.L.T. This significant reduction in training time is due to the shorter sequence lengths enabled by our compact factorized representation, making the model both efficient and effective for video generation tasks. Additionally, our factorization and recombination layers are lightweight, adding minimal overhead to the encoding and decoding processes. Additional timing details are included in the appendix.
4.1.3 Training and architecture details
To ensure a fair comparison against W.A.L.T. [20], we adopt a similar training setup and network architecture. The outer layers of our autoencoder architecture mirror the MAGVIT-v2 design to facilitate consistent benchmarking and reliable performance evaluations. In line with recent approaches in latent diffusion models, we simplify the autoencoder by removing the quantizer, working directly with continuous latent representations. To obtain the planes from the latent feature volume, we use average pooling as the factorization operations $\Lambda_{h}$ , $\Lambda_{w}$ and $\Lambda_{t}$. To recompose the planes into feature volume, we use the concatenation operation as $\mathrm{Combine}$. The autoencoder is trained with a combination of objectives, including an L2 reconstruction loss, a perceptual loss, and an adversarial loss, to ensure high-quality latent representations that preserve both fine details and overall structure. The L2 loss encourages accurate pixel-level reconstructions, while the perceptual loss focuses on retaining high-level semantic features, making the representation more robust. Additionally, the adversarial loss sharpens the output, helping the autoencoder produce visually realistic results [15].
To train the diffusion model we use a network architecture identical to W.A.L.T. We use a self-conditioning [8] rate of $0.9$, AdaLN-LoRA [20] with $r=2$ as the conditioning mechanism and zero terminal SNR [33] to avoid mismatch between training and inference arising from non-zero signal-to-noise ratio at the final time in noise schedules. We additionally use query-key normalization in the transformer to stabilize training. Our model is trained with a batch size of $256$ using an Adam optimizer with a base learning rate of $5\times 10^{-4}$ with a linear warmup and cosine decay.
@@ 4.2 Video interpolation
The video interpolation task evaluates our model’s ability to generate intermediate frames between given keyframes, testing both its temporal coherence and smoothness across sequences. Video interpolation is essential for applications requiring fluid motion reconstruction, such as frame-rate upsampling and video inpainting. In this experiment, we leverage our factorized latent representation to generate the spatio-temporal planes, $\mathbf{P}_{xt}$ and $\mathbf{P}_{yt}$, conditioned on the spatial planes, $\mathbf{P}_{xy}^{1}$ and $\mathbf{P}_{xy}^{2}$. As discussed in Sec. 3.1, the spatial-planes are obtained using a frozen image tokenizer.
The image tokenizer used in our model is trained on WebLI [9]. We train our video tokenizer, with spatial planes obtained from the image tokenizer, on an internal dataset similar to ViDiM, at a resolution of $256\times 256$ with 9 frames. We train the diffusion model on this same internal dataset, specifically curated for video tasks. The training and architectural setup closely follow the details outlined in Section 4.1.3. To evaluate our model’s interpolation capability, we test on the DAVIS-7 and UCF-7 datasets, as proposed in ViDiM [28]. These datasets consist of 400 videos, each containing 9 frames, and feature scenes with significant and often ambiguous motion.
As shown in Section 4.1.1, we summarize the quantitative results and compare them against various prior works using reconstruction-based metrics such as PSNR, SSIM, and LPIPS. Although these metrics are commonly used, they can penalize alternative, yet plausible, interpolations. To address this limitation, we also report FVD on the entire video, providing a more holistic evaluation of interpolation quality. We observe that our method performs comparably to VIDIM on the reconstruction metrics. Notably, unlike VIDIM—a diffusion-based baseline requiring a two-stage process with an initial base model followed by a super-resolution step—our model achieves $256\times 256$ resolution video generation in a single stage, making it both simpler and more efficient. We present qualitative results on two DAVIS scenes in Figure 4 to illustrate the effectiveness of our approach in video interpolation. Our method demonstrates comparable quality to the state-of-the-art VIDIM model while producing noticeably sharper details than other methods. These results emphasize the strength of our factorized representation in preserving fine textures and achieving high-fidelity frame generation in complex video scenes.
@@ 4.3 Ablation studies
To validate the effectiveness of our design choices, we perform ablations on the factorization and combine method as well as number of planes.
@@ 4.3.1 Factorization
We experiment with two variations of the factorization operation to optimize model performance. In the first variation, we apply mean pooling for $\Lambda_{h}$, $\Lambda_{w}$, and $\Lambda_{t}$, effectively reducing the dimensions while retaining essential features. In the second variation, we use a linear projection to map the feature dimension to $1$ along the targeted axis. When evaluating both models on the UCF class-conditional task Table 3, our results reveal that the mean pooling approach outperforms the linear projection variant, highlighting its effectiveness in balancing simplicity and performance.
@@
Table 3: Ablation: factorization method. We report FVD and Inception scores on the class-conditional task for the UCF-101 [44] dataset, comparing two factorization approaches: mean pooling and linear projection.
@@ 4.3.2 Combine
We also experiment with different variations of the combine operation to assess their impact on performance. Specifically, we test two approaches, concatenation and summation:
$\mathbf{V}(x,y,t)$ $=[\mathbf{f}^{1}_{xy}\lvert\lvert\mathbf{f}^{2}_{xy}\lvert\lvert\mathbf{f}_{yt}\lvert\lvert\mathbf{f}^{1}_{xt}]$
$\mathbf{V}(x,y,t)$ $=\mathbf{f}^{1}_{xy}+\mathbf{f}^{2}_{xy}+\mathbf{f}_{yt}+\mathbf{f}^{1}_{xt}$
We evaluate these variations on the UCF class-conditional task and present the results in Table 4. We find that concat yields a better performance as compared to summing the features of from the planes for volume creation. This improvement likely stems from concatenation’s ability to retain more distinct feature information from each plane.
@@
Table 4: Ablation: Combine method. We report FVD and Inception scores on the class-conditional task for the UCF-101 [44] dataset comparing performance across two different combination methods.
[…]
Four-plane TML]E3F6FF38 TML]E3F6FF91.13
Tri-plane 52 90.46
4.3.3 Triplane representations
A common approach to reduce dimensionality in volumetric data is the tri-plane representation [7]. While this representation can be effective for class-conditional generation, it is inherently limited for tasks like frame prediction and video interpolation due to the mixing of temporal information within the xy-plane. PVDM [55] employs this three-plane structure, achieving an FVD of $343$ and an Inception score of 74.40 on the UCF-101 dataset at a resolution of $256\times 256$. Our model achieves a significantly better FVD of $104$ and Inception score of $89.96$ on the same task. PVDM however used a 2D UNet based architecture compared to a transformer diffusion architecture in our model. Thus, to accurately assess the impact of the four-plane representation we perform an experiment where we substitute the four-plane represenation with tri-plane. We report the results in Table 5. Comparatively, the four-plane representation achieves better performance while also providing additional flexibility in its ability to be applied towards frame-conditional tasks in a straightforward manner.
@@ 5 Conclusion
Acknowledgements
We would like to thank Thomas Kipf for helpful discussions and feedback.
Appendix A Implementation details
A.1 Video autoencoder
To incorporate image data into the training of the video autoencoder, we adopt an image pretraining strategy commonly employed in prior works [51, 53, 20]. Specifically, we first train an image autoencoder using 2D convolutional layers. The trained weights are then used to initialize the video autoencoder. Following the approach in MAGVITv2 [53], which shares a similar architecture with our model, we inflate the 2D weights to 3D by initializing the 3D filters to zero and assigning the last slice of the 3D filter to the corresponding 2D filter weights. This method ensures a smooth transition from image-based training to video-based learning, leveraging the pre-trained image representations effectively.
A.2 Denoiser
  • Class-conditional generation: The denoiser is trained for $74,000$ iterations with a batch size of $256$. The input sequence has a length of $672$, comprising two spatial planes with a resolution of $16\times 16$ each and two spatio-temporal planes with a resolution of $5\times 16$ each.
A.3 Diffusion
Appendix B Timing details
@@
Appendix C Qualitative results
We provide frame interpolation results in frameprediction-video folder. Although our frame prediction model achieves a slightly higher FVD, it still generates plausible and visually coherent extrapolations.. We also provide additional results for interpolation on the DAVIS dataset in the interpolation-videos folder.
Appendix D Limitations
Our model was trained on $128\times 128$ resolution videos with 17 frames for the class-conditional generation and frame prediction. Our models performance on longer-duration videos remains unexplored. Similarly, the interpolation model was trained in a single stage. While it achieves competitive results, its quality could potentially be improved by scaling up the dataset, increasing the model size, and incorporating an upsampling stage.

### v2

@@ 4.1 Autoencoder reconstruction
Class Conditional Generation (FVD $\downarrow$) Frame Prediction (FVD $\downarrow$) Params Steps
UCF-101 (128x128) UCF-101 (256x256) Kinetics-600 (128x128)
Video Diffusion [23] - - 16.2 1.1B 256
RIN  [26] - - 10.7 411M 1000
Phenaki [51] - - 36.4 227M 1024
MAGVIT [56] 76 - 9.9 306M 48
PVDM [60] - 399.4 - - 400
MAGVIT-v2 [58] 58 - 4.3 307M 24
HVDM [30] - 303.1 - 63M 100
WALT [19] 46 - TML]B3E5FC3.3 313M 50
WALT* TML]E3F6FF39 TML]E3F6FF84.68 TML]E3F6FF5.7 214M 50
4Plane (Ours) TML]B3E5FC38 TML]B3E5FC58.27 8.6 214M 50
Table 2: Class-conditional generation on UCF and frame prediction on Kinetics-600. WALT* represents our re-training and re-evaluation of the WALT baseline. UCF-128 and UCF-256 refer to exeperiments at 128 x 128 and 256 x 256 resolutios respectively. Our method achieves competitive performance with WALT on the UCF-128 task and performs slightly lower on the K600 frame prediction task, showcasing efficient performance across both datasets. On UCF-256 our model outperforms the baselines.
4.2 Class-conditional generation
We evaluate the four-plane factorized latent space in a variety of generative settings, starting with class-conditional generation. Our evaluation is on the same two autoencoders described in the previous section, trained on K600 for 17 frame video reconstruction, at $128\times 128$ and $256\times 256$ resolutions. For generation, we train a transformer-based diffusion model to generate factorized latent embeddings, following Section 3.2. The diffusion model is trained on the UCF-101 dataset [47], which comprises 9,537 videos spanning 101 action categories, offering a diverse set of motion dynamics. To evaluate the quality of generated videos, we use the Fréchet Video Distance (FVD) [48] as our primary metric. FVD measures the similarity between the distributions of generated and real videos, assessing both spatial realism and temporal coherence.
Figure 3: Class-conditional generation results on the UCF dataset. We show every other frame of the 17-frame generated videos from the $128\times 128$ models. The temporal continuity and overall frame quality of our factorized model is comparable to the volumetric W.A.L.T. generations.
@@ 4.2 Class-conditional generation
4.2.1 Diffusion training details
For a fair comparison, we use a network architecture identical to W.A.L.T. Despite the architecture similarity, W.A.L.T. requires $1.6\times 10^{12}$ FLOPs whereas our model only uses $8.5\times 10^{11}$ due to shorter sequence lengths . We use a self-conditioning [8] rate of $0.9$, AdaLN-LoRA [19] with $r=2$ as the conditioning mechanism and zero terminal SNR [36] to avoid mismatch between training and inference arising from non-zero signal-to-noise ratio at the final time in noise schedules. We additionally use query-key normalization in the transformer to stabilize training. Our model is trained with a batch size of $256$ using an Adam optimizer with a base learning rate of $5\times 10^{-4}$ with a linear warmup and cosine decay.
@@ 4.2.2 Analysis
At a resolution of $128\times 128$, our factorized model outperforms most prior works and performs comparably to W.A.L.T. [19] (see  Tab. 2). Furthermore, our model achieves a higher Inception Score of 92.21, compared to 90.95 reported by W.A.L.T. Additionally, the smaller sequence length for our method makes diffusion training and inference nearly 2x faster compared to W.A.L.T. – under identical training architecture and computational resources, our model processes each training iteration in just 380 ms compared to 750 milliseconds for W.A.L.T. We show a detailed timing analysis in the appendix. A comparison of qualitative results are provided in Fig. 3.
At $256\times 256$ resolution, our factorization enables a diffusion model that outperforms prior baselines, despite the significant compression. Recall that the factorized sequence length for our $256\times 256$ resolution model is the same as for $128\times 128$, resulting in 4x more compression. Nonetheless, the generation quality decays relatively much less. Moreover, the comparison to WALT* indicates that a shortened sequence that retains critical spatiotemporal information can in fact reduce the modeling burden on the denoiser network, thereby improving generation quality. We show qualitative comparison in the appendix.
Compared to the tri-plane factorization approaches PVDM [60] and HVDM [30], the four-plane factorization yields substantial improvements, validating the critical design choices of our model. An additional tri-plane ablation is provided in the appendix. Another concern is that the tri-plane structure introduces information mixing in the spatial planes, making the representation unsuitable for different tasks like frame prediction, which we evaluate next.
@@
4.3 Future frame prediction
For the future frame predicition task, we reuse the autoencoder trained for class-conditional generation (Section 4.2). The denoiser network architecture and training procedure remain consistent with the details provided in Sec. 4.2.1, with the key difference that it is trained on K600 to align with the benchmark setting.. The diffusion model uses the first spatial plane $\mathbf{P}_{xy}^{1}$ as a conditioning sequence and learns to generate the remaining three planes $\mathbf{P}_{xy}^{2}$, $\mathbf{P}_{xt}$, and $\mathbf{P}_{yt}$. By leveraging a causal encoder, this setup mirrors W.A.L.T.’s frame prediction approach, where the model conditions on two latent frames. The frame prediction results in  Tab. 2 indicate our model outperforms most prior works and is comparable to WALT* while being significantly faster (see timing analysis in appendix).
@@ 4.4 Video interpolation
Video interpolation techniques generate intermediate frames between given keyframes, and are essential for applications requiring fluid motion reconstruction, such as frame-rate upsampling and video inpainting. In this experiment, we leverage our factorized latent representation to train the diffusion model to generate the spatio-temporal plane latents, $\mathbf{P}_{xt}$ and $\mathbf{P}_{yt}$, conditioned on the spatial plane latents, $\mathbf{P}_{xy}^{1}$ and $\mathbf{P}_{xy}^{2}$.
We train our video autoencoder and diffusion model on an internal dataset to encode and generate $256\times 256$ resolution videos with 9 frames. The training and architectural setup closely follow the details outlined in Section 4.2.1. We test on the DAVIS-7 and UCF-7 datasets, as proposed in VIDIM [27]. These datasets consist of 400 videos, each containing 9 frames, and feature scenes with significant and often ambiguous motion.
Section 4.2 shows the evaluations using reconstruction-based metrics such as PSNR, SSIM, and LPIPS. Although these metrics are commonly used, they can penalize alternative, yet plausible, interpolations. To address this limitation, we also report FVD on the entire video, providing a more holistic evaluation of interpolation quality. We observe that our method performs comparably to VIDIM on the reconstruction metrics. Notably, unlike VIDIM—a diffusion-based baseline requiring a two-stage process with an initial base model followed by a super-resolution step—our model achieves $256\times 256$ resolution video generation in a single stage, making it both simpler and more efficient. We present qualitative results on two DAVIS scenes in Figure 4 to illustrate the effectiveness of our approach in video interpolation. Our method demonstrates comparable quality to the state-of-the-art VIDIM model while producing noticeably sharper details than other methods. These results emphasize the strength of our factorized representation in preserving fine textures and achieving high-fidelity frame generation in complex scenes.
@@ 4.5 Ablation studies
To validate our design choices, we conduct ablations on both the factorization and combine methods across the class-conditional and frame prediction tasks, reporting FVD and Inception scores (IS) for the former, and FVD for the latter. We omit IS for frame prediction as it primarily evaluates classifiability and diversity making it unsuitable for the prediction task. See the appendix for additional ablation experiments.
@@ 4.5.1 Factorization
We explore two variations of the factorization operation (Sec. 3.1.1). The first approach applies mean pooling (MP) along $\Lambda_{h}$, $\Lambda_{w}$, and $\Lambda_{t}$, effectively reducing dimensionality while preserving essential features. The second employs a learned linear projection (LP) that maps the channel dimesion to $1$ along the targeted axis. Both methods perform comparably for frame prediction (Tab. 4, K600). However, for class-conditional generation (Tab. 4, UCF-101), MP significantly outperforms LP. We attribute this gap to the autoencoder’s limited generalization, as it is trained on K600. While both methods achieve similar reconstruction results on the K600 test set, with FVD scores of 7.8 (MP) and 8.0 (LP), their performance diverges on UCF, with FVD of 29.5 (MP) versus 37.1 (LP).
@@
K600 Mean Pooling 8.6 -
Linear Projection 8.9 -
Table 4: Factorization method. We contrast mean pooling with linear projection for factorizing the volumetric latents.
@@ 4.5.2 Combine
We also assess different choices for the combine operation (Sec. 3.1.2). Specifically, we test two approaches, concatenation defined as $\mathbf{V}(x,y,t)=[\mathbf{f}^{1}_{xy}\lvert\lvert\mathbf{f}^{2}_{xy}\lvert\lvert\mathbf{f}_{yt}\lvert\lvert\mathbf{f}^{1}_{xt}]$, and summation given by $\mathbf{V}(x,y,t)=\mathbf{f}^{1}_{xy}+\mathbf{f}^{2}_{xy}+\mathbf{f}_{yt}+\mathbf{f}^{1}_{xt}$. Table 5 shows that concatenation yields better performance across both the class-conditional generation and frame prediction task. This improvement likely stems from its ability to retain more distinct feature information from each plane.
@@
K600 Concat 8.6 -
Sum 27 -
Table 5: Combine method. We contrast concatenation with summation for recomposing the volume from the factorized latents.
@@ 5 Conclusion
Appendix A Frames vs Reconstruction Quality
We evaluate the performance of the four-plane factorized representation as the number of video frames increases. Specifically, we test videos with 17, 21, and 25 frames, and report reconstruction performance in Table 6. We observe that increasing the number of frames from 17 to 25 has only a minor impact on reconstruction quality.
[…]
17 27.11 0.82 0.051
21 26.95 0.82 0.051
25 26.51 0.81 0.052
Table 6: Video reconstruction. We report reconstruction metric of 4Plane tokenizer for various video lengths.
Appendix B Longer Video Generation
To demonstrate our model’s performance on longer sequences, we extend the class-conditional generation experiments to $36$ and $56$ frames, with results presented in Tab. 7. For $36$ frames, our method achieves comparable performance to WALT while running 5×5$\times$5 × faster. In the $56$-frame setup, WALT exceeds memory limits due to the increased sequence length, whereas our approach remains efficient.
FVD (ms/step)
36 frame 56 frames
WALT 164 (1350) OOM
4Plane (Ours) 171 (270) 564 (378)
Table 7: Longer Video Generataion. We report the FVD and training time per step (ms / step) for videos with 36 and 56 frame.
Appendix C Timing details
@@
During inference on a $128\times 128$ video with 17 frames, the 4Plane representation takes 0.17 seconds per video, while the volumetric representation takes 0.40 seconds per video, measured on a $2\times 2$ TPU v5e.
@@
Appendix D Triplane Ablation
[…]
Four-plane 38 91.13
Tri-plane 52 90.46
To assess the impact of the four-plane representation we perform an experiment where we substitute the four-plane represenation with tri-plane. We report the results in Table 8. We only conduct this ablation on the class-conditional task as the frame prediction task cannot be achieved with three plane representation. The superior performance of the four-plane representation over the tri-plane approach can be attributed to its increased capacity for capturing spatial information. By incorporating two spatial planes rather than a single one, the four-plane factorization preserves a more comprehensive set of spatial features, reducing information loss while also providing additional flexibility in its ability to be applied towards frame-conditional tasks in a straightforward manner.
Appendix E Implementation details
E.1 Video autoencoder
To incorporate image data into the training of the video autoencoder, we adopt an image pretraining strategy commonly employed in prior works [56, 58, 19]. Specifically, we first train an image autoencoder using 2D convolutional layers. The trained weights are then used to initialize the video autoencoder. Following the approach in MAGVITv2 [58], which shares a similar architecture with our model, we inflate the 2D weights to 3D by initializing the 3D filters to zero and assigning the last slice of the 3D filter to the corresponding 2D filter weights. This method ensures a smooth transition from image-based training to video-based learning, leveraging the pre-trained image representations effectively. For the $128\times 128$ experiments, the tokenizer consists of $4$ residual blocks in both the encoder and decoder, with $2$ temporal downsampling layers and $3$ spatial downsampling layers. At $256\times 256$, we increase the capacity to $5$ residual blocks, maintaining $2$ temporal downsampling layers while expanding to $4$ spatial downsampling layers. To trained the autoencoder we use a combination of objectives, including an L2 reconstruction loss, a perceptual loss, and an adversarial loss, to ensure high-quality latent representations that preserve both fine details and overall structure. For the VAE experiments in Section 4.1, we add an additional KL loss with a weight of $10^{-6}$
E.2 Denoiser
  • Class-conditional generation: The denoiser is trained for $74,000$ iterations with a batch size of $256$. For the $128\times 128$ resolution experiments, the input sequence has a length of $672$, comprising two spatial planes with a resolution of $16\times 16$ each and two spatio-temporal planes with a resolution of $5\times 16$ each. For the $256\times 256$ resolution experiments the dimension of the planes and thus the sequence length remains the same due to the additional temporal downsampling.
E.3 Diffusion
Appendix F Joint Image Video Training
While we have not experimented with it, our four-plane representation can be applied in the joint image-video training setting using the following strategy. When encoding a single image with our 4Plane encoder, the output consists of two identical spatial planes and two spatio-temporal vectors. One of the redundant spatial planes can be discarded, allowing us to use the remaining spatial plane along with the spatio-temporal vectors to train the denoiser network. This results in a slight increase in sequence length compared to the volumetric baseline. For example, given a latent grid of size $1\times 16\times 16$, the sequence length increases from 256 (volumetric) to 288 (4Plane), due to the inclusion of two additional spatio-temporal vectors of size 16 each. This strategy allows for using the same autoencoder and denoiser network for images and videos.